# Human4DiT: Free-view Human Video Generation with 4D Diffusion Transformer

### v1

@@ 4 Network Structure
Table 1: The collected multi-dimensional training dataset.
4.2 Control Modules
4.2.1 Camera Control Module
@@
$\begin{split}\mathbf{y}_{c}=(\sin(2^{0}\pi\mathbf{r}_{1}),\cos(2^{0}\pi\mathbf{r}_{1}),...,\\ \sin(2^{L-1}\pi\mathbf{r}_{1}),\cos(2^{L-1}\pi\mathbf{r}_{1}))\end{split}$ (4)
@@
$\mathbf{x}^{v^{\prime}}_{t}=\mathbf{x}^{v}_{t}+f_{c}(\mathbf{y}_{c})$ (5)
@@
4.2.2 Temporal Embedding Module
@@
$\mathbf{x}^{m^{\prime}}_{t}=\mathbf{x}^{m}_{t}+f_{m}(\mathbf{y}_{m})$ (7)
4.2.3 SMPL Control Module
To inject the SMPL parameters into the network, we first obtain the SMPL mesh vertex positions from the given shape and pose parameters $\mathbf{p}_{i}=LBS(\theta_{i},\beta_{i})$ of each frame $i$. We then render these vertices $\mathbf{p}_{i}$ into normal maps $\mathbf{M}_{n}(i,v)$ using camera parameters $\mathbf{c}_{v}$, as normals provide an effective representation of the 3D human. Since the camera parameters $\mathbf{c}_{v}$ have already been injected into the network, we decouple the SMPL information from the rendered normal maps by multiplying it with the inverse of the camera rotation matrix $\mathbf{r}_{v}^{-1}$. We then use a UNet $E_{p}$ to extract features $\mathbf{y}_{n}$ from these SMPL rendered normal maps $\mathbf{M}_{n}(i,v)$. These features $\mathbf{y}_{n}$ are eventually added to the input $\mathbf{x}_{t}$ of the 4D Diffusion Transformer before being fed into the network. Through this approach, the control condition is injected into the network in a pixel-aligned manner, ensuring consistent and accurate human motion video generation.
4.2.4 Human Identity Reference Module
During human video generation, we extract the human identity from a reference image and inject it into the network. We employ two ways to maintain identity consistency. First, similar to the SMPL injection, we use a UNet $E_{id}$ to extract human features $\mathbf{y}_{id}$ from the reference image. These features are then added to the input of the 4D diffusion transformer $\mathbf{x}_{t}$, allowing the network to better capture the detailed identity characteristics from the reference image. Additionally, we extract an image embedding $\mathbf{y}_{e}$ using CLIP and inject it into each transformer block via cross-attention after the self-attention mechanism:
$\begin{split}\mathbf{Q},\mathbf{K},\mathbf{V}&=f^{\mathbf{Q}}(\mathbf{\hat{x}}_{t}),f^{\mathbf{K}}(\mathbf{y}_{e}),f^{\mathbf{V}}(\mathbf{y}_{e}),\\ \mathbf{\tilde{x}}_{t}&=\text{softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}})\mathbf{V}\end{split}$ (8)
@@ 4.2.4 Human Identity Reference Module
Figure 3: Qualitative comparison on monocular video.
5 Dataset and Training Strategy
To train this 4D transformer model, we collected a multi-dimensional dataset and devised a multi-dimensional training strategy that fully leverages all available data modalities. As shown in Tab. 1, our multi-dimensional dataset comprises human images, videos, multi-view videos, 3D data, as well as a small amount of 4D data spanning different viewpoints and time steps.
@@ 5 Dataset and Training Strategy
This multi-dimensional training approach allows us to effectively leverage diverse data sources, with each modality contributing to different components of the model. The unified 4D transformer architecture seamlessly integrates information from all modalities during training, enabling coherent free-viewpoint human video generation.
@@ 15       $x_{t-1}\leftarrow\mu_{\theta}+\sigma_{t}\epsilon$
6 Efficient Spatial-Temporal Sampling
Our 4D diffusion transformer is capable of generating multi-view human motion videos. However, it cannot directly generate free-viewpoint videos where both viewpoint and time vary simultaneously. This limitation arises from computational memory constraints, where the total number of frames $M$ in the network input is bounded by the product of the temporal window size $M_{T}$ and the view window size $M_{V}$. To enable generation of long free-view videos, we propose an efficient and spatio-temporally consistent sampling method. During the denoising process, this method employs two strategies. The first strategy treats the free-view video as a monocular long video sequence, maximizing the temporal window $M_{T}^{1}$ to ensure long-term temporal consistency. The second strategy regards the free-view video as a collection of multi-view short video clips, using a larger viewpoint window $M_{V}^{2}$ with a smaller temporal window $M_{T}^{2}$, focusing on maintaining consistency across viewpoints. During denoising, the noise predictions from these two strategies are combined using respective weightings $\lambda_{1},\lambda_{2}$. The specific sampling algorithm is outlined in Alg. 1.
This sampling approach combines the advantages of both strategies. Globally, it leverages the view transformer to maintain consistency across large viewpoint separations. For smaller inter-frame motions between adjacent viewpoints, the temporal transformer ensures coherence. Ultimately, this method achieves spatio-temporally consistent generation of long free-viewpoint human motion videos under limited input window constraints.
Figure 4: Qualitative comparison multi-view video.
Figure 5: Qualitative comparison 3D static video.
Figure 6: Qualitative comparison free-view video.
7 Experiment
In this section, to validate the capabilities of our 4D diffusion transformer, we conducted comprehensive comparisons against current state-of-the-art human video generation methods including Disco [37], MagicAnimate [41], AnimateAnyone [7], and Champ [50] on monocular video, multi-view video, 3D static video, and free-view video generation.
@@ 7 Experiment
Table 2: Quantitative comparison on monocular video.
@@
Disco 18.86 0.796 0.293 646.2
MagicAnimate 19.30 0.845 0.232 517.3
AnimateAnyone 19.87 0.858 0.216 472.8
Champ 20.15 0.886 0.203 442.4
Ours (temporal) 21.16 0.907 0.190 374.2
Ours 22.40 0.920 0.159 296.53
Table 3: Quantitative comparison on multi-view video.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$
Disco 17.13 0.882 0.209 451.6
MagicAnimate 19.28 0.906 0.159 356.4
AnimateAnyone 20.53 0.922 0.111 285.4
Champ 21.11 0.927 0.106 204.3
Ours (temporal) 22.58 0.941 0.058 165.2
Ours 23.37 0.962 0.045 110.0
Table 4: Quantitative comparison on 3D static video.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$
Disco 19.98 0.872 0.169 559.7
MagicAnimate 21.74 0.920 0.135 418.4
AnimateAnyone 22.01 0.912 0.131 410.3
Champ 23.35 0.922 0.110 347.6
Ours (temporal) 24.74 0.936 0.085 274.1
Ours 25.02 0.947 0.062 234.8
Table 5: Quantitative comparison on Free-view video.
7.1 Implementation Details
@@
Our method is based on the latent diffusion model that utilizes the VAE from Stable Diffusion XL. Our 4D diffusion transformer contains 10 4D transformer blocks, totaling 30 transformer layers. Our U-Net compresses the input by a factor of 8, with each token having 1280 channels, and the text embedding and image identity embedding are also mapped to 1280 channels. During training, the video data is resized to 768x768 resolution, and for each GPU, the input consists of 24 frames. When training on image data, we use a batch size of 24, while for monocular video training, we use a batch size of 1 with a video length of 24. For multi-view video, 3D video, and free-view video training, we use a batch size of 1 with a video length of 6 and 4 views. We employed 24 A100 GPUs, and the total training time was 14 days.
7.2 Comparisons
@@
Comparisons on Monocular Video. For monocular video, we randomly select 200 videos from the Diffuhuman-video dataset as our test set for comparison, with the first frame of each video serving as the reference image. During inference, our method employs a temporal window of 24 frames, with an overlap of 6 frames between consecutive windows. For Disco [37], MagicAnimate [41], and Champ [50], we use their official open-source code. As for AnimateAnyone [7], we employ the opensource implementation from MooreThreads https://github.com/MooreThreads/Moore-AnimateAnyone. Quantitative comparisons are presented in Tab. 2, where our method demonstrates a clear numerical advantage, significantly outperforming other approaches. This highlights the superiority of our 4D diffusion transformer over U-Net-based methods. We have also conducted qualitative comparisons, with results shown in Fig. 3. Our method generates more natural dynamic effects with fewer deformation and jitter artifacts compared to other methods, indicating the 4D transformer’s stronger capability in establishing spatio-temporal consistency than U-Net-based approaches. Please refer to our submitted project webpage for dynamic video effects.
@@
Comparisons on Multi-view Video. For multi-view settings, we select 25 multi-view groups from the AIST and DNA-rendering datasets respectively as our test set for comparison, with each group containing videos captured from 4 different viewpoints. We use the first frame of the frontal view as the reference image. During inference, our method employs the spatial-temporal sampling strategy, with a spatial window size of 4 views and a temporal window of 6 frames for spatial sampling. For temporal sampling, the window size is 24 frames. The blending weights for spatial and temporal $\lambda_{1},\lambda_{2}$ are set to 0.5 and 0.5, respectively. For Disco, MagicAnimate, Champ, and AnimateAnyone, we treat each view’s video as a monocular video and perform inference separately. Quantitative comparisons are presented in Tab. 3, where we masked out the backgrounds for metric computation since inferring other views’ backgrounds from a single view is an ill-posed problem. Our method outperforms others, demonstrating the 4D diffusion transformer’s superior ability to establish stronger cross-view correlations compared to U-Net-based approaches. We have also conducted qualitative comparisons, with results shown in Fig. 4. Our method generates spatio-temporally consistent multi-view videos without exhibiting multi-face artifacts. Please refer to the project webpage for dynamic video effects.
@@
Comparisons on 3D Static Video. For 3D setting, we select 100 3D scans from THUman2.0 and render them into 3D static videos as our test set for comparison. We use the frontal view as the reference image. During inference, our method employs the same spatial-temporal sampling strategy as the multi-view setting. Quantitative comparisons are presented in Tab. 4. We mask out the backgrounds for a fair comparison since other methods tend to generate noisy backgrounds. Our method outperforms others, demonstrating the 4D diffusion transformer’s ability to learn physical 3D viewpoint changes. We have also conducted qualitative comparisons, with results shown in Fig. 5. Our method generates spatially consistent free-view videos compared to other approaches. Please refer to our submitted project webpage for 3D static video effects.
@@
Comparisons on Free-view Video. For free-view video evaluation, we select 10 4D scans from the Human4DiT dataset, each performing 5 different motions, and render them with different camera trajectories to create 50 test videos. We use the frontal view as the reference image. During inference, our method employs the same spatial-temporal sampling strategy as the multi-view setting. Quantitative comparisons are presented in Tab. 5, where we mask out the backgrounds for the evaluation. The results demonstrate our 4D diffusion transformer’s ability to handle both viewpoint and human motion changes in a 4D dynamic scenario. We have also conducted qualitative comparisons, with results shown in Fig. 6. Our method generates spatio-temporally consistent free-viewpoint dynamic videos compared to other approaches. Please refer to our submitted project webpage for free-view video effects.
7.3 Ablation Study

### v2

@@ 4. Network Structure
Table 1. The collected multi-dimensional training dataset.
4.1. 4D Diffusion Transformer
Our 4D diffusion transformer performs denoising on a randomly initialized noisy multi-view latent human video $\mathbf{z}_{t}$. After multiple denoising steps, it produces the denoised output $\mathbf{z}_{0}$, which is eventually decoded by a VAE to generate the video. We first utilize a CNNs-based encoder $\mathbf{E}_{\mathbf{z}}$ to extract latent tokens. Then the input could be regarded as a set of tokens in 5D ($\mathbf{z}_{t}\in\mathbb{R}^{V\times T\times H\times W\times C}$), where $C$ represents the feature dimension of each token, $V$ is the viewpoint dimension, $T$ is the time dimension, and $H$ and $W$ are the height and width, respectively. To establish correlations across viewpoints, time, and 2D spatial dimensions, we propose a novel cascaded diffusion transformer architecture. It could decompose the complex 4D attention into three types of attention including image, temporal and view attention, which efficiently learns the 4D correspondences and decreases the memory usage.
First, we feed $z_{t}$ into a 2D image transformer block to capture spatial self-attention within each frame:
(1) $\begin{split}\mathbf{z}_{t}^{s}&=\text{rearrange}(\mathbf{z}_{t},(V\times T,H\times W,C)),\\ \mathbf{Q}_{s},\mathbf{K}_{s},\mathbf{V}_{s}&=f_{s}^{\mathbf{Q}}(\mathbf{z}_{t}^{s}),f_{s}^{\mathbf{K}}(\mathbf{z}_{t}^{s}),f_{s}^{\mathbf{V}}(\mathbf{z}_{t}^{s}),\\ \mathbf{\hat{z}}_{t}^{s}&=\text{softmax}(\frac{\mathbf{Q}_{s}\mathbf{K}_{s}^{T}}{\sqrt{d_{k}}})\mathbf{V}_{s},\end{split}$
where $f_{s}^{\mathbf{Q}},f_{s}^{\mathbf{K}},f_{s}^{\mathbf{V}}$ are the linear layers in the transformer block. The rearrange operation reshapes the tensor $\mathbf{z}_{t}$ into a 3D tensor $\mathbf{z}_{t}^{s}\in R^{(V\times T)\times(H\times W)\times C}$, where the first dimension $V\times T$ is treated as the batch size. The actual attention operation is then performed on the last two dimensions $H\times W$ and $C$, which correspond to the 2D spatial dimensions of the images. After the 2D image transformer, we further model correspondences across different viewpoints through the view transformer block:
(2) $\begin{split}\mathbf{z}_{t}^{v}&=\text{rearrange}(\mathbf{\hat{z}}_{t}^{s},(T,V\times H\times W,C)),\\ \mathbf{Q}_{v},\mathbf{K}_{v},\mathbf{V}_{v}&=f_{v}^{\mathbf{Q}}(\mathbf{z}_{t}^{v}),f_{v}^{\mathbf{K}}(\mathbf{z}_{t}^{v}),f_{v}^{\mathbf{V}}(\mathbf{z}_{t}^{v}),\\ \mathbf{\hat{z}}_{t}^{v}&=\text{softmax}(\frac{\mathbf{Q}_{v}\mathbf{K}_{v}^{T}}{\sqrt{d_{k}}})\mathbf{V}_{v},\end{split}$
In the view transformer, due to the substantial variations across different viewpoints, especially between frontal, side, and back views, we perform attention jointly across the viewpoint and the 2D spatial dimensions. This attention allows for better modeling of the global correlations across the entire frame. It also makes the view transformer the most computationally and memory-intensive component, limiting the maximum viewpoint window size. We will address this constraint during the inference process, as discussed later. After the view transformer, we finally employ a temporal transformer to capture temporal correlations across time steps.
(3) $\begin{split}\mathbf{z}_{t}^{m}&=\text{rearrange}(\mathbf{\hat{z}}_{t}^{v},(V\times H\times W,T,C)),\\ \mathbf{Q}_{m},\mathbf{K}_{m},\mathbf{V}_{m}&=f_{m}^{\mathbf{Q}}(\mathbf{x}_{t}^{m}),f_{m}^{\mathbf{K}}(\mathbf{z}_{t}^{m}),f_{m}^{\mathbf{V}}(\mathbf{z}_{t}^{m}),\\ \mathbf{\hat{z}}_{t}^{m}&=\text{softmax}(\frac{\mathbf{Q}_{m}\mathbf{K}_{m}^{T}}{\sqrt{d_{k}}})\mathbf{V}_{m},\end{split}$
The three transformer blocks (2D image, view, and temporal) are interconnected to form a single 4D transformer block. Our complete 4D diffusion transformer architecture is composed of 10 cascaded 4D transformer blocks. Through this cascaded multi-level attention scheme, our approach significantly reduces the computational overhead while improving the training efficiency and effectively ensuring spatio-temporal consistency in the generation process.
4.2. Control Modules
4.2.1. Camera Control Module
@@
(4) $\mathbf{y}_{c}=(\sin(2^{0}\pi\mathbf{r}_{1}),\cos(2^{0}\pi\mathbf{r}_{1}),...,(\sin(2^{L-1}\pi\mathbf{r}_{1}),\cos(2^{L-1}\pi\mathbf{r}_{1}))$
@@
(5) $\mathbf{z}^{v^{\prime}}_{t}=\mathbf{z}^{v}_{t}+f_{c}(\mathbf{y}_{c})$
@@
4.2.2. Temporal Embedding Module
@@
(7) $\mathbf{z}^{m^{\prime}}_{t}=\mathbf{z}^{m}_{t}+f_{m}(\mathbf{y}_{m})$
4.2.3. SMPL Control Module
To inject the SMPL parameters into the network, we first obtain the SMPL mesh vertex positions from the given shape and pose parameters $\mathbf{p}_{i}=LBS(\theta_{i},\beta_{i})$ of each frame $i$. We then render these vertices $\mathbf{p}_{i}$ into normal maps $\mathbf{M}_{n}(i,v)$ using camera parameters $\mathbf{c}_{v}$, as normals provide an effective representation of the 3D human. Since the camera parameters $\mathbf{c}_{v}$ have already been injected into the network, we decouple the SMPL information from the rendered normal maps by multiplying it with the inverse of the camera rotation matrix $\mathbf{r}_{v}^{-1}$. We then use a CNNs-based Encoder $\mathbf{E}_{p}$ to extract features $\mathbf{y}_{n}$ from these SMPL rendered normal maps $\mathbf{M}_{n}(i,v)$. These features $\mathbf{y}_{n}$ are eventually added to the input $\mathbf{x}_{t}$ of the 4D Diffusion Transformer before being fed into the network. Through this approach, the control condition is injected into the network in a pixel-aligned manner, ensuring consistent and accurate human motion video generation.
4.2.4. Human Identity Reference Module
During human video generation, we extract the human identity from a reference image and inject it into the network. We employ two ways to maintain identity consistency. First, similar to the SMPL injection, we use a CNNs-based encoder $E_{id}$ to extract human features $\mathbf{y}_{id}$ from the reference image. These features are then added to the input of the 4D diffusion transformer $\mathbf{x}_{t}$, allowing the network to better capture the detailed identity characteristics from the reference image. Additionally, we extract an image embedding $\mathbf{y}_{e}$ using CLIP and inject it into each transformer block via cross-attention after the self-attention mechanism:
(8) $\begin{split}\mathbf{Q},\mathbf{K},\mathbf{V}&=f^{\mathbf{Q}}(\mathbf{\hat{z}}_{t}),f^{\mathbf{K}}(\mathbf{y}_{e}),f^{\mathbf{V}}(\mathbf{y}_{e}),\\ \mathbf{\tilde{z}}_{t}&=\text{softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}})\mathbf{V}\end{split}$
@@ 4.2.4. Human Identity Reference Module
5. Dataset and Training Strategy
To train this 4D transformer model, we collected a multi-dimensional dataset and devised a multi-dimensional training strategy that fully leverages all available data modalities. As shown in Tab. 1, in addition to utilizing publicly available human datasets, we collect a new multi-dimensional dataset comprises human images, videos, multi-view videos, 3D data, as well as a small amount of 4D data spanning different viewpoints and time steps. It contains 5k 3D human scans capture by camera rigs, 10k videos from YouTube/Bilibili, and 100 animatable human models from artists. Some samples are presented in Fig. 3. We adopt two methods to estimate SMPL for our dataset: For multi-view and 3D/4D data, we detect 2D key points and optimize SMPLs to align with the corresponding images (Zhang et al., 2021). For monocular videos, we directly use Humans-in-4D (Goel et al., 2023) to estimate image-aligned SMPLs and cameras.
@@ 5. Dataset and Training Strategy
This multi-dimensional training approach allows us to effectively leverage diverse data sources, with each modality contributing to different components of the model. The unified 4D transformer architecture seamlessly integrates information from all modalities during training, enabling coherent 360-degree human video generation.
@@ 15       $x_{t-1}\leftarrow\mu_{\theta}+\sigma_{t}\epsilon$
6. Efficient Spatial-Temporal Sampling
Our 4D diffusion transformer is capable of generating multi-view human motion videos. However, it cannot directly generate 360-degree videos where both viewpoint and time vary simultaneously. This limitation arises from computational memory constraints, where the total number of frames $M$ in the network input is bounded by the product of the temporal window size $M_{T}$ and the view window size $M_{V}$. To enable generation of long 360-degree videos, we propose an efficient and spatio-temporally consistent sampling method. During the denoising process, this method employs two strategies. The first strategy treats the 360-degree video as a monocular long video sequence, maximizing the temporal window $M_{T}^{1}$ to ensure long-term temporal consistency. The second strategy regards the 360-degree video as a collection of multi-view short video clips, using a larger viewpoint window $M_{V}^{2}$ with a smaller temporal window $M_{T}^{2}$, focusing on maintaining consistency across viewpoints. During denoising, the noise predictions from these two strategies are combined using respective weightings $\lambda_{1},\lambda_{2}$. The specific sampling algorithm is outlined in Alg. 1.
This sampling approach combines the advantages of both strategies. Globally, it leverages the view transformer to maintain consistency across large viewpoint separations. For smaller inter-frame motions between adjacent viewpoints, the temporal transformer ensures coherence. Ultimately, this method achieves spatio-temporally consistent generation of long 360-degree human motion videos under limited input window constraints.
7. Experiment
In this section, to validate the capabilities of our 4D diffusion transformer, we conducted comprehensive comparisons against current state-of-the-art human video generation methods including Disco (Wang et al., 2023), MagicAnimate (Xu et al., 2023), AnimateAnyone (Hu et al., 2023), and Champ (Zhu et al., 2024) on monocular video, multi-view video, 3D static video, and 360-degree video generation.
Figure 4. Qualitative comparison on monocular video.
@@ 7. Experiment
AnimateAnyone* 24.27 0.862 0.148 341.0
Champ* 24.93 0.870 0.139 307.6
Ours (small) 25.45 0.877 0.121 276.1
Ours (image) 24.15 0.858 0.147 491.2
@@
Table 2. Quantitative comparison on monocular video. * indicates that we fine-tuned the model on our dataset. Ours (small) is the small version of our 4D transformer model with fewer layers. Ours (image) indicates that our model only has image transformer blocks.
@@
MV 3D 4D MV 3D 4D MV 3D 4D MV 3D 4D
Disco 18.86 17.13 19.98 0.796 0.882 0.872 0.293 0.209 0.169 646.2 451.6 559.7
MagicAnimate 19.30 19.28 21.74 0.845 0.906 0.920 0.232 0.159 0.135 517.3 356.4 418.4
AnimateAnyone 19.87 20.53 22.01 0.858 0.922 0.912 0.216 0.111 0.131 472.8 285.4 410.3
Champ 20.15 21.11 23.35 0.886 0.927 0.922 0.203 0.106 0.110 442.4 204.3 347.6
AnimateAnyone* 20.78 21.92 24.06 0.904 0.937 0.931 0.193 0.068 0.089 413.0 194.9 301.2
Champ* 20.90 22.18 24.34 0.905 0.940 0.935 0.185 0.060 0.083 397.3 186.3 286.1
Ours (small) 22.02 22.94 24.56 0.911 0.952 0.935 0.166 0.053 0.069 344.3 129.7 256.5
Ours (image) 20.66 21.41 24.15 0.899 0.931 0.928 0.197 0.072 0.088 563.1 319.0 435.4
Ours (image+temporal) 21.16 22.58 24.74 0.907 0.941 0.936 0.190 0.058 0.085 374.2 165.2 274.1
Ours 22.40 23.37 25.02 0.920 0.962 0.947 0.159 0.045 0.062 296.53 110.0 234.8
Table 3. Quantitative comparison on multi-view (MV), 3D static (3D), and 360-degree (4D) videos. * indicates that we fine-tuned the model on our dataset. Ours (small) is the small version of our 4D transformer model with fewer layers. Ours (image) indicates that our model only has image transformer blocks. Ours (image+temporal) indicates our model without the view transformer blocks.
7.1. Implementation Details
7.1.1. Network Architecture
@@
Our method is based on the latent diffusion model that utilizes the VAE from Stable Diffusion XL. Our 4D diffusion transformer contains 10 4D transformer blocks, totaling 30 transformer layers. Our CNNs-based encoder including $\mathbf{E}_{p},\mathbf{E}_{id}$ compresses the input by a factor of 8, with each token having 1280 channels, and the text embedding and image identity embedding are also mapped to 1280 channels. The latent encoder $\mathbf{E}_{\mathbf{z}}$ further compresses the input by a factor of 2. During training, the video data is resized to 768x768 resolution, and for each GPU, the input consists of 24 frames. When training on image data, we use a batch size of 24, while for monocular video training, we use a batch size of 1 with a video length of 24. For multi-view video, 3D video, and 360-degree video training, we use a batch size of 1 with a video length of 6 and 4 views.
7.1.2. Training Details
We employed 24 A100 GPUs, and the total training time was 14 days. We train our model with a learning rate 1e-5, first on the image dataset for 50k iterations, then on the full dataset for 150k iterations. We use ColossalAI’s HybridAdam optimizer with Zero2 strategy and zero weight decay to optimize parameters. We also clip the gradient in the range of [-1, 1].
7.2. Comparisons
7.2.1. Comparisons on Monocular Video.
@@ 7.2. Comparisons
For monocular video, we randomly select 200 videos from the Human4DiT-Video dataset as our test set for comparison, with the first frame of each video serving as the reference image. During inference, our method employs a temporal window of 24 frames, with an overlap of 6 frames between consecutive windows. For Disco (Wang et al., 2023), MagicAnimate (Xu et al., 2023), and Champ (Zhu et al., 2024), we use their official open-source code. As for AnimateAnyone (Hu et al., 2023), we employ the opensource implementation from MooreThreads  [^1] https://github.com/MooreThreads/Moore-AnimateAnyone . To ensure a more fair comparison, we fine-tuned AnimateAnyone and Champ on our Human4DiT dataset for 10k iterations. Results from these fine-tuned models are denoted with an asterisk (*) in our evaluations. Quantitative comparisons are presented in Tab. 2, where our method demonstrates a clear numerical advantage, significantly outperforming other approaches. This highlights the superiority of our 4D diffusion transformer over CNNs-based methods. We have also conducted qualitative comparisons, with results shown in Fig. 4. Our method generates more natural dynamic effects with fewer deformation and jitter artifacts compared to other methods, indicating the 4D transformer’s stronger capability in establishing spatial-temporal consistency than U-Net-based approaches. Please refer to our submitted project webpage for dynamic video effects.
Figure 5. Ablation study of spatial-temporal sampling.
7.2.2. Comparisons on Multi-view Video.
@@
For multi-view settings, we select 25 multi-view groups from the AIST and DNA-rendering datasets respectively as our test set for comparison, with each group containing videos captured from 4 different viewpoints. We use the first frame of the frontal view as the reference image. During inference, our method employs the spatial-temporal sampling strategy, with a spatial window size of 4 views and a temporal window of 6 frames for spatial sampling. For temporal sampling, the window size is 24 frames. The blending weights for spatial and temporal $\lambda_{1},\lambda_{2}$ are set to 0.5 and 0.5, respectively. For Disco, MagicAnimate, Champ, and AnimateAnyone, we treat each view’s video as a monocular video and perform inference separately. Quantitative comparisons are presented in Tab. 3, where we masked out the backgrounds for metric computation since inferring other views’ backgrounds from a single view is an ill-posed problem. Our method outperforms others, demonstrating the 4D diffusion transformer’s superior ability to establish stronger cross-view correlations compared to U-Net-based approaches. We have also conducted qualitative comparisons, with results shown in Fig. 8. Our method generates spatio-temporally consistent multi-view videos without exhibiting multi-face artifacts. Please refer to the project webpage for dynamic video effects.
7.2.3. Comparisons on 3D Static Video.
@@
For 3D setting, we select 100 3D scans from THUman2.0 and render them into 3D static videos as our test set for comparison. We use the frontal view as the reference image. During inference, our method employs the same spatial-temporal sampling strategy as the multi-view setting. Quantitative comparisons are presented in Tab. 3. We mask out the backgrounds for a fair comparison since other methods tend to generate noisy backgrounds. Our method outperforms others, demonstrating the 4D diffusion transformer’s ability to learn physical 3D viewpoint changes. We have also conducted qualitative comparisons, with results shown in Fig. 9. Our method generates spatially consistent 360-degree videos compared to other approaches. Please refer to our submitted project webpage for 3D static video effects.
7.2.4. Comparisons on 360-degree Video.
@@
For 360-degree video evaluation, we select 10 4D scans from the Human4DiT dataset, each performing 5 different motions, and render them with different camera trajectories to create 50 test videos. We use the frontal view as the reference image. During inference, our method employs the same spatial-temporal sampling strategy as the multi-view setting. Quantitative comparisons are presented in Tab. 3, where we mask out the backgrounds for the evaluation. The results demonstrate our 4D diffusion transformer’s ability to handle both viewpoint and human motion changes in a 4D dynamic scenario. We have also conducted qualitative comparisons, with results shown in Fig. 10. Our method generates spatio-temporally consistent 360-degree dynamic videos compared to other approaches. Please refer to our submitted project webpage for 360-degree video effects.
7.2.5. Comparisons on model size.
In addition to comparing video generation quality, we conducted a comprehensive analysis of model sizes across various approaches, as illustrated in Tab 4. We also train a smaller version of Human4DiT ”Ours (small)” with 6 4D transformer blocks. Our Human4DiT model in its small setting exhibits comparable parameter counts to AnimateAnyone and Champ. However, due to the higher computational complexity of attention operations in the DiT architecture compared to convolutional operations, our model does not demonstrate an advantage in terms of inference speed. Nonetheless, as presented by our previous experimental comparisons, our model achieves superior performance within fewer parameters, thus validating the efficiency of our proposed Human4DiT architecture.
Figure 6. Ablation study of temporal transformer. When using only an image transformer without a temporal transformer, temporal consistency could not be guaranteed, resulting in artifacts such as discontinuous arm generation (see top images). After introducing the temporal transformer, our method can produce continuous and natural human motion (see bottom images).
Figure 7. Failure cases. Our method produces distortions and artifacts in cases where the facial region is small (see left images). Additionally, our approach struggles to generate natural finger structures (see right images).
Method Parameters T-size Resolution FPS
AnimateAnyone 2.07B 24 512x512 0.769
Champ 2.04B 24 768x768 0.657
Ours (small) 1.83B 24 768x768 0.605
Ours 3.10B 24 768x768 0.443
Table 4. Performance comparisons of different video models. T-size is the temporal window size. FPS is the average number of generated frames every second with 25 inference steps.
7.3. Ablation Study
7.3.1. Temporal Transformer.
First, we validate our temporal transformer blocks and conduct an ablation study by removing the temporal transformers. The quantitative and qualitative results ”Ours (image)” are presented in Tab. 3 and Fig. 6. Observations indicate that the incorporation of temporal transformers significantly enhances generation performance. Notably, even in the absence of these components, our method demonstrates comparable performance to CNN-based approaches.
7.3.2. View Transformer